

********************************************************************
UARS WEEKLY STATUS REPORT - 27 JUL 2001
********************************************************************
>>>> UARS Operations to Cease on Sep. 30, 2001 <<<<
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
The cancellation of UARS has left SUSIM without needed reprocessing
to incorporate recent calibrations. It is a major setback. We very
much want this decision reversed.
SUSIM has completed its daily solar scans this past week during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
The PEM team reports that the PEM:: node computer has failed and has
been taken off line. PEM:: appears to have hardware problems that take the
system down every few hours. Automatic transfer NRT tests to PEM2K from
both CDHF2 and CDHFB have succeed and the NRT files have been verified against
those captured by PEM, which confirm the NRT process is correct. Thus, PEM
confirms NRT data from CDHFB. Thanks to Steve Bove for his work assisting
PEM with this process.
PEM/AXIS was activated in activated in full power mode on July 3 and
deactivated on July 19. At the current time, AXIS is not generating energy
deposition. SwRI is looking into this. Tests of AXIS production code on the
rehost computer are continuing.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE was powered on beginning-of-day today. All engineering parameters
are within nominal operating region. Gain and bore sight calibration tests
will resume and continue until the yaw turn off.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
On 25 July 2001 the MPG attended the UARS All-Hands
meeting discussing the recent decision to cease UARS
operations as of September 30 2001.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 14 July 2001 (Orbit 53776, GMT Day 195) through 20 July 2001
(Orbit 53895, GMT Day 201).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 198/1944 53848 Power Off
PEM 200/1007 53872 AXIS Off
The beta angle ranged between the angles of 63.0 degrees to 77.4 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 78 degrees on 20 July 2001 (DOY 201).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.2 degrees C. Additional battery performance data for each
battery on 13 July 2001 and 20 July 2001 are provided in Appendix D.
The clock error ranged from Ð11.8 to 12.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 06
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 04
¥ SUSIM = 01
¥ WINDII = 01
¥ PEM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 04
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/186 show the observatory data loss to be 302 hours, 44
minutes, 57.624 seconds. This shows a loss of 04 minutes and 37.544 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 04 minutes and 37.544 seconds occurred over five (5) days:
· 36.528 seconds on Day 196 due to a Data Hit.
· 2 minutes and 10.176 seconds on Day 197 due to Data Hits
· 59.392 seconds on Day 198 due to a Data Hit
· 1 minute and 50.448 seconds on Day 200 due to a late AOS TTR#14170
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
An instrument anomaly was experienced on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT conducted analysis and command tests to determine the cause of the
anomaly. Analysis shows that the anomaly was caused by failed circuitry in the
ACRIM electronics box. The FOT is awaiting a future plan of action from the
ACRIM team.
The software test team, in conjunction with the FOT, has completed a large
portion of the testing for the new front end processor. The new TAC still has
yet to undergo DSN and realtime TDRS testing, but has completed simulations of
receipt of telemetry, commanding, and NCC/GCMR testing. The DSN test scheduled
for 18 July 2001 was cancelled and has yet to be rescheduled due to problems
with the simulator.
UARSFOT will also begin supporting testing of the rehosted UARS ground system.
A prototype system has been built in accordance with the UARS documentation and
operational database. The demo of the prototype system will be held on 02
August 2001.
UARS attitude has been adjusted over the course of the last few yaw cycles by
loading new alignment matrices for the TAM. Good alignments have been found
for both directions of flight and the performance of the TAM should remain
stable and accurate.
MEETINGS
The UARS FOT attended a Flight Software CCB on Monday, July 16,2001.
The UARS FOT attended a 3 TDRS Meeting on Tuesday, July 17, 2001.
The UARS FOT attended a Safehold Working Group Meeting on Thursday, July 19,
2001.
The UARS FOT attended a TAC RS DR Meeting on Friday, July 20, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 August 2001 (DOY 220). This
maneuver will be a Forward to Reverse Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 20 July 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
53827 197/1006 01-458 OBC/CLK Clk Adj 20.6 to 41.2 msec/day
53848 198/1944 01-397 WINDII Power OFF
53847 198/1847 01-459 SUSIM Update sunset value
53858 199/1150 01-460 OBC/CLK Clk Adj 41.2 to 20.6 msec/day
53872 200/1007 01-461 PEM AXIS 1/2 POWER OFF
53882 201/0246 01-462 PEM (Z/H) FROM DAY ONLY TO FULLTIME OPS
53885 201/0745 01-463 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
53887 201/1008 01-464 OBC/CLK Clk Adj 0.00 to 41.2 msec/day
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 20 July 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
HRPTSBARTMP Y-LO PTS Barrel Temp 53815-20
HRPTSBARTMP Y-LO PTS Barrel Temp 53821-24
HRPTSBARTMP Y-LO PTS Barrel Temp 53855-57
HRPTSBARTMP Y-LO PTS Barrel Temp 53866-70
MLSCANATTMP Y-HI MLS Scanner Actuator Temp 53869-71
MLSCANENTMP Y-HI MLS Scan Encoder Temp 53869-71
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 20 July 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53827 197/1008 Clk Adj 20.6 to 41.2 msec/day
53858 199/1155 Clk Adj 41.2 to 20.6 msec/day
53885 201/0752 Clk Adj 20.6 to 0.00 msec/day
53887 201/1011 Clk Adj 0.00 to 41.2 msec/day
___________________________________________________________________________
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 20 July 2001
13 July 2001 - GMT Day 194 Beta = 59.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-33.6 +3.87/+2.60 +22.4/-14.4 26.4 1.030 18.5
3 +22.4/-22.4 +2.92/+1.66 +21.6/-14.4 26.4 1.020 18.0
20 July 2001 - GMT Day 201 Beta = 77.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -1.79/-3.05 +0.0/+0.0 12.2 0.000 0.0
2 +11.2/-89.6 +2.60/+1.34 +14.4/-15.2 31.5 0.096 13.3
3 +16.8/-44.8 +1.97/+1.03 +14.0/-15.6 31.5 0.095 13.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APENDIX E
UARS ANOMALY REPORTS
Week ending 20 July 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-041 FOT 53824 197/0503 14170 Generic Late Acquisition
01-041 FOT 53846 198/1705 14170 Generic Late Acquisition
01-041 FOT 53862 199/1848 14170 Generic Late Acquisition
01-041 FOT 53879 200/2207 14170 Generic Late Acquisitio
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-041 FOT see list July 01 14170 Generic Late Acquisition
SEVNTFW01/29